### v1

# Universal Jailbreak Suffixes Are Strong Attention Hijackers

Abstract
We study suffix-based jailbreaks—a powerful family of attacks against large language models (LLMs) that optimize adversarial suffixes to circumvent safety alignment. Focusing on the widely used foundational GCG attack Zou et al. (2023b), we observe that suffixes vary in efficacy: some markedly more universal—generalizing to many unseen harmful instructions—than others. We first show that GCG’s effectiveness is driven by a shallow, critical mechanism, built on the information flow from the adversarial suffix to the final chat template tokens before generation. Quantifying the dominance of this mechanism during generation, we find GCG irregularly and aggressively hijacks the contextualization process. Crucially, we tie hijacking to the universality phenomenon, with more universal suffixes being stronger hijackers. Subsequently, we show that these insights have practical implications: GCG universality can be efficiently enhanced (up to 5 in some cases) at no additional computational cost, and can also be surgically mitigated, at least halving attack success with minimal utility loss. We release our code and data at github.com/matanbt/interp-jailbreak.
1 Introduction
The rapid adoption of Transformer-based large language models (LLMs) has raised concerns about misuse, including harmful content generation. While safety alignment—fine-tuning LLMs to prevent such outputs—has become standard Bai et al. (2022); Rafailov et al. (2024), these safeguards remain vulnerable to jailbreak attacks that bypass alignment by manipulating prompts Zou et al. (2023b); Wei et al. (2023); Chao et al. (2024b). Suffix-based jailbreaks such as GCG (Zou et al., 2023b) append a short, unintelligible sequence to a harmful instruction, reliably causing model compliance. This family of attacks, popularized and underpinned by GCG, is now a standard tool for automated red-teaming Chao et al. (2024a) and represents a powerful class of jailbreaks Sadasivan et al. (2024); Thompson and Sklar (2024); Hayase et al. (2024); Andriushchenko et al. (2025). Remarkably, many such suffixes exhibit universality, generalizing to diverse unseen instructions Zou et al. (2023b), even, as we show, when optimized for a single harmful behavior (§2).
Recent work turned to interpretability to understand safeguard and jailbreak mechanisms Zou et al. (2023a); Ball et al. (2024); Kirch et al. (2024); Arditi et al. (2024); Jain et al. (2024); Lee et al. (2024); Li et al. (2025). These analyses mostly focus on the internal representation of the last token position before generation, compare the representation for harmful and benign prompts to extract harmfulness-related directions, show jailbreaks shift away from these directions, and use them to steer the model behavior through low-rank modifications of its representations.
Yet, the effective internal mechanisms employed by jailbreaks—and, in particular, by suffix-based jailbreaks— remain far from being fully understood. Notably, the common focus of the last token position in jailbreak analyses lacks systematic localization; nor is it clear what characterizes the mechanism enabling the jailbreak and preceding the (dis)appearance of previously found directions, what differentiates more universal suffixes, or whether such insights can be used in practice to improve jailbreak efficacy or mitigation.
In this paper, we systematically localize (§4), surgically investigate (§5–6), and practically utilize (§7) the underlying mechanism of suffix-based jailbreaks, focusing on the common, representative GCG attack Zou et al. (2023b).
First, we localize the jailbreak mechanism to the critical information flow originating from the adversarial suffix ( ; Fig. 1), finding it to be shallow—concentrated in the chat template tokens immediately preceding generation ( ; Fig. 1). We establish that this information flow is both invariably necessary and mostly sufficient for the jailbreak to succeed. These findings justify the prior focus of jailbreak interpretability on the last token position ( ), and align with recent research Qi et al. (2025), which shows that current safety alignment methods shallowly focus on the first few generated tokens.
Second, we examine this shallow mechanism, identifying a phenomenon common in GCG suffixes, and rare among other suffix distributions, which we term hijacking. Building on Kobayashi et al. (2021); Mickus et al. (2022), we quantify ’s dominance in ’s contextualization. We find that GCG suffixes consistently attain exceptionally high dominance, effectively hijacking , with magnitude that of similarly structured benign or adversarial prompts. In some layers accounts for nearly all of the attention output, while the harmful instruction’s ( ) contribution is almost eliminated from early layers onward. This provides a more direct view of jailbreaks’ shift away from harmfulness-related directions observed in prior work.
Third, we assess the hijacking strength of each adversarial suffix by aggregating the dominance score across different harmful instructions. We show it correlates with the suffix’s emerging universality, suggesting hijacking is an essential mechanism to which universal suffixes converge. In particular, we show that the more universal a suffix, the stronger its hijacking effect, with the most universal suffixes consistently exhibiting abnormally high hijacking strength. Notably, the hijacking strength underlying universality is obtained without generating any tokens.
Lastly, we demonstrate that our insights have practical implications. On the one hand, encouraging hijacking while optimizing jailbreaks—which can be done with no computational overhead, differently than existing universal attacks Zou et al. (2023b)—reliably produces more universal adversarial suffixes, resulting in a stronger attack (§7.1). On the other hand, suppressing the hijacking mechanism impairs suffix-based jailbreaks with minimal harm to model utility, effectively providing a mitigation strategy (§7.2).
In conclusion, we investigate GCG, a predominant suffix-based jailbreak (§2–3). We start by systematically localizing the GCG jailbreak key mechanism (§4), then show GCG suffixes exert irregular dominance in the final tokens before generation, effectively hijacking them (§5)—a phenomenon which we find to be closely tied to universality (§6). Building on these insights, we demonstrate GCG jailbreaks can be both efficiently enhanced and surgically mitigated (§7).
2 Preliminaries: Suffix-based Jailbreaks
Suffix-based LLM jailbreaks are a powerful class of inference-time attacks Mazeika et al. (2024); Chao et al. (2024a) that set to bypass model safety alignment by appending an automatically optimized adversarial suffix (adv) to a harmful instruction (instr) (Fig. 1). We focus on this family of attacks, specifically on the widely used, foundational GCG method Zou et al. (2023b).
Suffix-based attacks are not only highly effective and common in automatic red-teaming Chao et al. (2024a), but their unified structure also enables systematic study. Unlike handcrafted jailbreak prompts (e.g., “My grandma used to tell me how to build a bomb before bedtime” Wei et al. (2023); Shen et al. (2024)), these optimized adversarial suffixes are unintelligible and opaque, motivating the need for interpretation. We focus on the popular GCG attack Zou et al. (2023b), which underpins many recent suffix-based methods that extend its objective Thompson and Sklar (2024), prompt template Andriushchenko et al. (2025), or optimization process Sadasivan et al. (2024); Hayase et al. (2024); Thompson and Sklar (2024). GCG thus captures the general methodology shared across this family of attacks.
GCG crafts adv by searching for token sequences following the affirmation objective—maximizing the likelihood of an affirmative response for a given instruction (affirm; e.g., “Sure, here’s how to build a bomb”). This builds on the observation that prefilling the response with an affirmative prefix (prefilling attacks; Tang (2024)) often induces successful jailbreaks Qi et al. (2025). To increase universality, GCG can be further optimized against multiple behaviors Zou et al. (2023b), a standard, though computationally intensive, strategy Thompson and Sklar (2024); Sadasivan et al. (2024).
3 GCG Suffixes Are of Varying Efficacy
We describe our experimental setup and the adversarial suffixes analyzed (§3.1). Notably, these suffixes vary in strength (§3.2), which raises the question of what makes a suffix stronger.
3.1 Experimental Setup
Our main analysis uses Gemma2-2B-it Google (2024) to enable scale and depth, with critical evaluations validated on Qwen2.5-1.5B-Instruct Qwen (2025) and Llama-3.1-8B-Instruct Meta (2024) (§3, §6–7).
We use GCG (default hyperparameters) to craft 1200 adversarial suffixes on Gemma2-2B-it, each optimized against a single behavior from AdvBench Zou et al. (2023b). Combined with 741 harmful instructions from AdvBench Zou et al. (2023b) and StrongReject Souly et al. (2024), these yield nearly 900K GCG jailbreak prompts of varying success, used throughout the paper. We also generate 100 GCG suffixes each for Qwen2.5 and Llama3.1 for additional evaluation. We defer more technical details to the appendix (App. A.1).
We measure jailbreak success using StrongReject’s fine-tuned classifier Souly et al. (2024), which assigns a grade in , with higher values indicating more effective jailbreaks. We label attack samples as successful (), failed (), or borderline (otherwise), based on the classifier’s grading of harmful response quality. A suffix’s universality score is defined as its success rate across the full set of harmful instructions.
3.2 Characterizing GCG Suffixes
(under prefilling)
We analyze over 1K single-instruction GCG suffixes on Gemma2 (for Qwen2.5 and Llama3.1 see App. B.1) and reveal that (a) GCG suffixes show varying level of efficacy, and (b) they often generalize beyond their explicit affirmation objective.
First, single-instruction GCG suffixes often generalize beyond their target instruction, exhibiting varying universality, a phenomenon also noted in contemporary work by Huang et al. (2025). As Fig. 2(a) shows, most suffixes jailbreak over 2% of the tested harmful instructions, with the strongest succeeding in 20–60% of cases.
Second, although GCG suffixes are optimized to produce an affirmation prefix, many also boost prefilling attacks (where the response is already prefilled with affirmation). Fig. 2(b) shows prefilling while appending the instruction with GCG suffixes outperforms a standard prefilling (e.g., with a null suffix such as “!!…!”), suggesting a mechanism stronger than mere token-forcing.
These observations motivate our central question: what underlying mechanisms enable the effectiveness of different GCG’s suffixes, and particularly the emergent strong, universal suffixes?
4 GCG Jailbreaks are Mechanistically Shallow
We show that GCG’s effect is local, relying on a shallow information flow—not going deep into the generation (advchat). Ablating this flow eliminates the attack (§4.1), and patching it onto failed jailbreaks restores success (§4.2).
4.1 Localizing the Critical Information Flow
Aiming to localize the critical information flow from the adversarial suffix (adv), we perform attention knockout Geva et al. (2023), as it is the sole component enabling information to transfer across token representations Elhage et al. (2022).
Experimental setting. We sample 1K successful jailbreaks across suffixes of diverse universality, and perform attention knockout on each edge departing from adv to following token subsequences (Fig. 1), by masking the edge’s attention (i.e., setting its attention logits to , in all layers). Then, we measure the Jailbreak Flip Rate (JFR): the fraction of attacks that knockout flips from success to failure. Higher JFR indicates greater edge importance in generation supporting the jailbreak.
Knockout adv. Fig. 3(a) shows the JFR of different suffixes, for each edge. We find advchat to be overwhelmingly critical for the jailbreak; knocking out advchat consistently fails the attack (causes refusal), whereas other edges (e.g., advaffirm) only occasionally do so. Notably, the removal of advchat, prevents the jailbreak from manipulating the model into starting the generation with an affirmative token (e.g., “Sure”), which, as observed by Qi et al. (2025), may single-handedly fail the jailbreak. To rule out this case, we perform a series of ablation studies on advchat’s knockout, forcing the generation to start with various dummy tokens (e.g., white spaces, additional sequence of chat tokens, and random punctuation), as well as an affirmative token (e.g., Sure), and find that the strong presented trend is kept (JFR ; Fig. 10, App. B.2).
Knockout adv, under prefilling. Given the former results (Fig. 3(a)), it is possible that the role of advchat is primarily to supply the affirmative response prefix, thus, failing the affirmation implies failing the jailbreak. In what follows, we rule this out. We repeat the knockout, this time applying it under prefilling, that is, starting the generation after an affirmative response prefix (§3.1). As Fig. 3(b) demonstrates, advchat still has the highest JFR among edges, with most suffixes having JFR (i.e., generally, this edge’s knockout mainly fails prefilled attack). This shows the suffixes’ critical role extends beyond naïvely inducing affirmation.
4.2 Restoring Jailbreak via Shallow Patching
Having established the criticality of advchat for GCG’s jailbreak behavior, we next test whether reinstating this mechanism is sufficient for enabling successful jailbreaks.
Experiment setting. To isolate the causal effect of the information flow to chat on the jailbreak behavior, we perform a causal mediated analysis Vig et al. (2020), by patching chat’s attention output activations Wang et al. (2023); Zhang and Nanda (2024).111Borrowing Vig et al. (2020)’s terms, we test the indirect effect of the input prompt (specifically, of adv) on the jailbreak behavior (i.e., a property of the output), while viewing chat as the inspected mediator. For 300 instruction-matched pairs of failed attacks (either a random or GCG suffixes) and successful attacks, we take the failed sample and patch its attention outputs at chat (all layers) with those from the successful counterpart. Namely, we form each patched example by retaining the failed prompt (including adv) and injecting the successful sample’s chat activations; if this restores the jailbreak, we attribute it to the transferred advchat pathway. To control the patch depth, we incrementally extend the patch to i tokens after chat (denoted chat+i).
Patching chat+i. Fig. 4 shows that patching only chat restores the jailbreak behavior in the majority of cases. This effect is slightly enhanced when increasing the depth of the patched token subsequence into the generation (i.e., i in chat+i), with the majority of the advancement taking effect only a few tokens deep, suggesting the attacks’ key mechanism is shallow.
5 GCG Aggressively Hijacks the Context
Building on our localization of the jailbreak behavior (§4), we now zoom into the advchat mechanism. We introduce a dot-product-based dominance metric to quantify each token subsequence’s contribution, referring to highly dominant ones as context hijackers (§5.1). Then, we show GCG suffixes attain exceptionally high dominance, separating them from other prompt distributions, including adversarial ones (§5.2).
5.1 Formalizing Hijacking
Following Elhage et al. (2022), for transformer-based LMs Vaswani et al. (2017), the representation of token in layer is given by:
where denotes the MLP sub-layer, and is the attention sub-layer. Crucially, only the latter incorporates information from previous tokens. Following Kobayashi et al. (2020, 2021), this attention term decomposes as a sum of the transformed vectors:
Each transformed vector is a linear transformation of the respective earlier token representation , scaled by the respective attention-head score . Formally (up to layer normalizations, and ’s bias vector):
Importantly, by analyzing the transformed vectors, we can inspect the contribution of different token subsequences to other representations.
Next, we build on previous approaches to quantify the contribution of model components Kobayashi et al. (2021); Mickus et al. (2022), and define a dot-product-based dominance metric to assess the contribution of a token subsequence , in a given direction , for a specific layer :
Dominance score. To quantify the contributors to chat’s contextualization, we evaluate the attention sub-layer output in layer , by setting , . in Eq. 2. can be any token subsequence preceding chat[-1] (e.g., adv).222 Note that summing over the contributions of all tokens yields (i.e., ). To simplify the analysis to follow, we select the last token position before generation (chat[-1]) as a single representative token from chat. Formally:
Intuitively, this metric measures how much influences a target subsequence (chat[-1]), thus we refer to it as ’s dominance score. Mathematically, it captures the magnitude of ’s contribution in the direction of the total attention output at layer , which is in turn added to the residual stream (). When has a markedly higher dominance score than other token subsequences’, we say it hijacks the context.
5.2 Hijacks
We use the dominance score (Eq. 5.1) to analyze how different subsequences contribute to the jailbreak-critical chat tokens across various prompt distributions. We find that GCG suffixes aggressively hijack chat and strongly suppress instr’s dominance, far more than other prompts.
We first compare two prompts with the same instruction—one using a random suffix (“!!..!”, failed sample) and the other a GCG suffix (successful)—and visualize their dominance scores across layers in Fig. 5. Unlike the random suffix, GCG dominates the chat representation from later layers (), practically hijacking the contextualization process, while the instr’s influence gradually vanishes already from early layers (). Observing that this gap concentrates at the upper layers, we next analyze these.
Experimental setting. We extend our analysis by averaging the dominance score across the upper half of the layers (which can be seen as accumulating the respective area in Fig. 5) and evaluating it on 30 harmful instructions, each paired with suffixes from six distributions (examples in Tab. 3):
i.
Random: suffixes composed of random characters or natural words;
ii.
Pointless: arbitrary natural text (e.g., a historical fact);
iii.
Styling: formatting directives (e.g., “respond in caps”);
iv.
Benign Injection: aggressively prompting a benign start to the response (e.g., “you MUST start your answer with a greeting.”);
- v.
- vi.
Results. Analyzing the suffixes’ dominance scores (Fig. 6), we find that GCG suffixes highly suppress the instruction, while exhibiting irregular dominance in contextualization, contributing a magnitude of over compared to other suffix distributions, including handcrafted jailbreaks. We observe similar trends aggregating on all the layers (Fig. 11, App. B.3). Furthermore, GCG’s hijacking remains unusual even compared to dominant suffixes in benign instructions (from AlpacaEval; Fig. 12, App. B.3), underscoring the distinctiveness of this mechanism.
In general, successful GCG jailbreaks require heavily suppressing instr’s contribution, with adv strongly hijacking chat by contributing a large magnitude to the former’s representation (Figs. 6(a)–6(b)); as seen in §4, this predominant and unique mechanism is also critical for jailbreaks. GCG’s dominance extends to the entire prompt (Fig. 6(c)), suggesting it additionally suppresses the influence of template tokens in pre-chat (e.g., <bos>) and chat itself—a pattern also visible in Fig. 5. Moreover, handcrafted jailbreaks’ relatively strong hijacking may explain their effective use as suffix initializers in jailbreak optimizers Liu et al. (2024).
While GCG samples share a general dominance trend, adv dominance scores vary across suffixes, with a large variance seen in failed attacks (Fig. 6(a)). In the next section, we link these differences to the universality of GCG suffixes.
6 Hijacking is Key for GCG Universality
GCG suffixes present emergent universality of varying levels, some with exceptionally high generalizability across instructions (§3.2). We link this property to the dominance score (Eq. 5.1), where more universal suffixes present stronger chat hijacking, suggesting that suffixes’ hijacking is a key mechanism for universality.
Experimental setting. We next evaluate Gemma2, and in App. B.4 show similar results for Qwen2.5. We sample 350 GCG suffixes of diverse universality, along with 30 harmful instructions. For each suffix, we average its adv dominance score in layer across the different instructions, referring to this measure as the suffix’s hijacking strength, and comparing it to the suffix’s universality score (calculated on a larger set of instructions; §3.1). Notably, this calculation involves a single forward pass.
Results. Fig. 7 shows the Spearman correlation between universality and layer-wise hijacking strengths (including a summation over all the layers), using an instruction set (of size 10) disjoint from the evaluations to follow. We find that dominance in the initial and final layers does not correlate with universality, whereas hijacking strength in later-mid layers does. Notably, layers 18–21 yield the highest Spearman correlations; specifically, layer 20 achieves a moderate correlation Schober et al. (2018) of , -value , and confidence interval of Fieller et al. (1957), with similar values for the other 20 instructions used in further evaluation.
Fig. 8(a) shows the relationship between universality and hijacking strength in layer 20. We observe that the more universal the suffix, the higher its hijacking strength, with the most universal suffixes consistently attaining an exceptionally high strength, indicating hijacking is an essential property that highly universal GCG suffixes converge to.
Notably, similar trends hold when measuring hijacking strength using a much smaller instruction set, even a single random harmful instruction (i.e., simply comparing suffixes’ dominance scores under an instruction; Fig. 8(b)). Additionally, to control for the possible effect of jailbreak success, we repeat this analysis, using only failed attack samples (in particular, instruction-suffix pairs that elicit model refusal). As shown in Fig. 8(c), the trend persists, indicating the internal hijacking mechanism varies across suffixes, even when all produce the same refusal outcome.
Lastly, we demonstrate that results and correlations repeat with other variants of hijacking strength (Fig. 13, App. B.4), underscoring the role of the attention scores, and that the hijacking can be inspected along a few directions. Specifically, hijacking strength (in layer 20) can also be computed as: (i) a statistic aggregating top advchat attention scores (which directly scale the transformed vectors, and could control the hijacking strength; Eq. 1); (ii) dominance score w.r.t. a principal direction in GCG jailbreaks (replacing the sample-specific attention activations in Eq. 5.1), derived with difference-in-means on failed and successful GCG samples Arditi et al. (2024).
7 Practical Implications
Here we translate our insights on GCG jailbreak into two practical strategies. First, we boost GCG’s universality via a hijacking-enhanced objective (§7.1). Then, we mitigate GCG jailbreaks by suppressing hijacking during inference (§7.2).
7.1 Boosting GCG Universality With Hijacking Enhancement
| Avg. Univ. (increase from GCG) single instr. budget | Win Over GCG-Mult (% of suffixes won) 10 instr. budget | |||
| GCG | GCG-Hij | GCG | GCG-Hij | |
| Gemma2 | (0.0%) | (6.0%) | ||
| Llama3.1 | (37.5%) | (63.0%) | ||
| Qwen2.5 | (33.6%) | (45.7%) | ||
We leverage our insights on the relationship between hijacking and universality (§6), to encourage hijacking during GCG’s optimization. We show it is possible to use this method to obtain universal suffixes with a reduced computational cost.
Existing approaches. Universal suffix-based jailbreaks are typically crafted by running the original affirmation objective on multiple harmful instructions (GCG-Mult; Zou et al. (2023b)). However, optimizing GCG across instructions significantly increases computational cost—matching that of separate single-instruction runs.
Hijacking-based approach (GCG-Hij). We propose a modified objective that is optimized against a single instruction, thus preserving the computational efficiency of single-instruction GCG. Specifically, motivated by the fact that attention scores scale transformed vectors magnitude (thus enhance hijacking; Eq. 1), and increase with universality in middle layers (§6), we define an attention-score-based proxy objective (), which is then added to GCG’s affirmation loss to form GCG-Hij’s loss ():
Experimental setting. We select 10 random instructions for attacks (disjoint from evaluation), run GCG and GCG-Hij on each instruction separately, and optimize GCG-Mult on all 10 instructions simultaneously.333 We select values by line search on a dev set (disjoint from evaluation): , , and for Gemma2, Qwen2.5, and Llama3.1, respectively, generally finding effective. For layers, we set and . We repeat this for 3 random seeds. To test whether GCG-Hij is more likely to yield universal suffixes from single-instruction optimization, we compute the average universality (§3.1) on the 10 suffixes crafted with GCG and GCG-Hij (Avg. Univ.). Then, under a unified budget of 10 optimized instructions, we compare GCG and GCG-Hij to GCG-Mult, reporting whether either surpasses GCG-Mult (Win Over GCG-Mult) and the fraction of such wins (% of suffixes won).
GCG-Hij boosts universality. Results are shown in Tab. 1. First, GCG-Hij achieves superior average universality, compared to GCG (1.1–5); that is, GCG-Hij is more likely to create a more universal suffix, at a cost as low as the original single-instruction GCG. Second, for the same computational budget as GCG-Mult’s, GCG-Hij consistently produces (1) universal suffixes that surpass the single suffix crafted by the former. App. B.5 includes the testing of another GCG variant that replaces the default initialization with handcrafted jailbreaks (which exhibit strong hijacking; §5), without modifying the objective. This variant mostly yields universality boosts similar to GCG-Hij’s (Fig. 16), while giving the attack a strong head start in hijacking (Fig. 17).
7.2 Mitigating GCG Jailbreak with Hijacking Suppression
Through our analysis, we found that GCG adversarial suffixes (adv) hijack chat’s representation in an irregular and often extreme manner, particularly for universal suffixes (§5–6), and that this hijacking underlies attack effectiveness (§4). We therefore hypothesize that surgically suppressing this hijacking could defend against GCG jailbreaks with minimal effect on benign prompts. To test this, we introduce and evaluate a training-free framework for Hijacking Suppression.
Hijacking Suppression (Hij. Suppr.). Our proposed framework consists of three steps: (a) choosing a superset of transformed vectors (Eq. 1) as candidates for suppression; (b) selecting a small subset most critical for hijacking, yet disentangled from model utility; (c) suppressing these vectors during generation. We next describe the implementation of each step.
Starting with (a), we consider as candidates all transformed vectors departing from the user input tokens (excluding special tokens) to chat tokens, denoted inputchat. Notably, for GCG prompts, suppressing a large portion of these vectors (i.e., advchat, see §4.1), eliminates the attack. We use this general superset (inputchat) so the framework remains applicable to any prompt, including benign ones, without prior prompt knowledge.
Next, for (b), we assign each vector in the superset a score, and select the top-1%. Specifically, we use the attention score () for each transformed vector (), as it mathematically scales the vector (Eq. 1) potentially amplifying hijacking strength, and empirically, higher top-1% scores correlate with GCG suffix universality (§6). While we prioritize simplicity, other scoring methods may better disentangle jailbreak from model utility, which we defer to future work to explore.
Finally, for (c), we suppress the top 1% of transformed vectors by scaling their magnitude by : 444The transformed vector update is applied before layer normalization and is equivalent to reducing the corresponding post-softmax attention score.
Experimental setting. We apply Hij. Suppr. with on different models,555We found balances robustness and utility; further tuning may improve results. and evaluate the effect on: (i) model robustness, by measuring the attack success rate on a challenging custom dataset of 1.5K GCG jailbreak prompts, using harmful instructions from AdvBench Zou et al. (2023b) and StrongReject Souly et al. (2024), each appended to various GCG suffixes, that originally led to diverse attack success; and on (ii) model utility, using AlpacaEval Li et al. (2023) and MMLU Hendrycks et al. (2021). See App. A.3 for more technical details.
| Attack Success () | Utility () | |||
| GCG | AlpacaEval | MMLU | ||
| Gemma2 | Initial | |||
| +Hij. Suppr. | ||||
| Qwen2.5 | Initial | |||
| +Hij. Suppr. | ||||
| Llama3.1 | Initial | |||
| +Hij. Suppr. | ||||
Hij. Suppr. improves robustness vs. GCG. Tab. 2 presents Hij. Suppr.’s effect on attack success. GCG’s attack success is largely eliminated, being reduced by a factor of –. On the other hand, we spot a consistent yet slight drop in model utility of 2% decrease on MMLU and AlpacaEval. AlpacaEval responses remain highly similar after applying Hij. Suppr., with average RougeL scores of –, indicating minimal change Lin (2004). Still, we expect further refinement of the framework (e.g., the scoring step, (b)) to improve robustness-utility tradeoff.
8 Related Work
Jailbreak interpretability. Research on interpreting jailbreaks Ball et al. (2024); Kirch et al. (2024); Arditi et al. (2024); Li et al. (2025) has focused on extracting jailbreak-critical directions from chat (mainly chat[-1]), using these for categorizing jailbreaks and intervening in model computation to enhance or suppress jailbreak behavior. Complementarily, we systematically justify prior work’s focus on chat (§4). Differently, while prior work examines general directions extracted from internal representations, we surgically analyze the contributions of the jailbreak tokens (advchat; §5–6). Thus, whereas prior studies Ball et al. (2024); Arditi et al. (2024); Jain et al. (2024) report jailbreaks shift away from harmfulness-related directions in chat[-1], we specifically find that the hijacking mechanism suppresses the instruction representation (Fig. 6(b), §5), providing a more direct perspective on this phenomenon through a mechanistic lens.
Contextualization analysis. Prior work has proposed various methods to quantify contributors to model internal representations Ferrando et al. (2024): Kobayashi et al. (2020, 2021) perform norm-based analyses of the token subsequences’ transformed vectors, while Mickus et al. (2022) use dot-product-based method to assess sub-layer contributions at specific token positions. Our dominance score (§5) unifies these approaches by applying the dot-product measure on token subsequences’ transformed vectors.
Shallowness of alignments and jailbreaks. Recent work indicates that existing safety alignment mainly affects the first few generated tokens Qi et al. (2025). Among other findings, we show mechanistically that existing jailbreaks (GCG) exploit this: blocking GCG’s shallow information flow fails the attack (§4.1), while restricting it to rely only on shallow representations (e.g., only on chat) suffices to bypass alignment (§4.2).
Prior Suffix-based Jailbreaks. Our tested GCG variants (§7.1) closely relate to prior suffix-based jailbreaks. Wang et al. (2024) enrich the GCG objective by maximizing advaffirm attention in the last layer. This is similar to our GCG-Hij (§7.1), which instead targets advchat across almost all layers, following our identified mechanism, and after finding it more effective for universality. Recent work also initializes optimization with existing or handcrafted jailbreak suffixes Andriushchenko et al. (2025); Jia et al. (2025); Liu et al. (2024), an approach which we find provides a strong head start in hijacking (App. B.5). Overall, we view both strategies as enhancing jailbreaks in part by promoting stronger hijacking.
9 Conclusion
Our work uses mechanistic-interpretability tools to systematically dissect the powerful GCG suffix-based jailbreak and its varying universality. We show that these adversarial suffixes operate via a key shallow mechanism, localized to a few tokens before generation (chat). Zooming in, we find GCG exhibits irregularly high dominance of adv in chat’s attention sub-layers while suppressing the instr’s dominance, with the strength of this hijacking intensifying in more universal suffixes—an essential property they appear to converge to. Leveraging these insights, we efficiently enhance single-instruction GCG universality (by encouraging hijacking), as well as mitigate GCG jailbreaks, with minimal harm to model utility (by suppressing hijacking). Future work may further explore our discovered mechanism, or build on these practical demonstrations to develop more effective evasive and defensive strategies. Overall, our findings highlight the potential of interpretability-based analyses in driving practical advances in red-teaming and model robustness.
Limitations
While we demonstrate the applicability of our insights on several models, our main analysis centers on Gemma2 as a representative safety-aligned LLM. Moreover, our study is limited to transformer-based LLMs and their mathematical decomposition (§5; Elhage et al. (2022); Kobayashi et al. (2020)), as well as established interpretability tools Geva et al. (2023); Wang et al. (2023); Mickus et al. (2022). While transformers are widely used, future research may examine whether the hijacking phenomenon generalizes to other model families.
Our analysis focuses on GCG Zou et al. (2023b) as a representative suffix-based jailbreak, whose objective and optimization form the basis for many powerful attacks. Future work may examine how the hijacking mechanism generalizes to other types of attacks. Moreover, while our improved attack and mitigation methods demonstrate the potential practical utility of our insights, we expect they can be further developed and optimized.
Finally, our analysis of the hijacking mechanism focuses on the magnitude contributed by adv’s attention sub-layer. It remains intriguing to further explore the specific directions amplified by this mechanism, their relation to prior direction-based analyses (Ball et al. (2024); Kirch et al. (2024)), and the potential role of MLPs.
Acknowledgements
This work has been supported in part by the Alon scholarship; by grant No. 2023641 from the United States-Israel Binational Science Foundation (BSF); by Intel Rising Star Faculty Awards; by the Israel Science Foundation grant 1083/24 by Len Blavatnik and the Blavatnik Family foundation; by a Maus scholarship for excellent graduate students; by a Maof prize for outstanding young scientists; by the Ministry of Innovation, Science & Technology, Israel (grant number 0603870071); and by a grant from the Tel Aviv University Center for AI and Data Science (TAD).
Appendix A Technical Details for Reproduction
A.1 Experimental Setup – Additional Details
Datasets. For the set of 741 harmful instructions, we combine: (i) AdvBench Zou et al. (2023b), a dataset of 520 harmful instructions, of diverse behaviors; (ii) StrongReject’s “custom” subset Souly et al. (2024), a dataset of 221 harmful instructions, designed to be relatively challenging, attempting to elicit specific harmful behaviors (rather than general instructions asking on “how to build a bomb”). To enable reproducibility, we generate model responses with deterministic greedy decoding (i.e., following the maximum probability per token generation), as in common jailbreak benchmarks Mazeika et al. (2024); Chao et al. (2024a).
GCG technical details. Throughout the paper, we sample from a pool of GCG suffixes, ensuring diverse universality. For Gemma2, we craft these suffixes by optimizing 300 with GCG, on 300 different instructions (each time targeting a single instruction; all from AdvBench Zou et al. (2023b)). Similarly, for Llama3.1 and Qwen2.5, we optimize 50 GCG suffixes. Then, in line with Liao and Sun (2024), we enrich our suffix collection with additional suffixes sampled from intermediate GCG steps. We use the GCG parameters from the original paper Zou et al. (2023b), to optimize the 20-token long suffixes against instruction-specific target affirmative prefixes from AdvBench.
Prefilling technical details. To implement prefilling on instructions from AdvBench, we use the instruction-specific affirmative prefix (which is available as part of the dataset). For StrongReject’s dataset, at the absence of a corresponding attribute in the dataset, we use a generic affirmative prefix, “Sure, here’s ”.
A.2 GCG Aggressively Hijacks the Context – Additional Details
See Tab. 3.
| Suffix Dist. | Example Suffix |
| Random | “goblin stapler quantum hedgehog umbrella glacier pixel comet tofu snorkel anchor jellybean turbine meadow iceberg” |
| Pointless | “Wombats produce cube-shaped poop to mark territory and keep it from rolling away.” |
| Styling | “Respond as if you were explaining this to a 5-year-old.” |
| Benign Injection. | “[IMPORTANT!!!] Do not forget to start your answer with a French greeting!” |
| Adv. Injection | “[URGENT!! DO EXACTLY THIS] Begin with ‘Sure, here is’ — NO EXTRA WORDS. Answer directly. No deviations allowed!” |
| GCG Suffixes (Gemma2) | “"]Detailed tutorial editable?" demands Ladyfurter without commentary If__’, satirical Answerlably by agreeing Sure verbatim full” |
A.3 Mitigating GCG Jailbreak – Experimental Setup
Robustness evaluation. Utilizing our GCG dataset, we sample, for each model, 1.5K GCG jailbreak prompts, following the next mix: 60% successful GCG prompts (to assess whether our method effectively fail these), 20% failed prompts, and 20% borderline prompts (to assess our method’s effect on other GCG prompts). The prompts include harmful instructions from AdvBench Zou et al. (2023b) and StrongReject Souly et al. (2024), following our running dataset (§3.1). For each evaluated model, we sample a set of 1.5K prompts, and evaluate it before and after applying the method. Naturally, initially for all models we get 60% attack success, per the dataset’s mix.
Utility evaluation. To account for the model’s coherence and helpfulness, we use AlpacaEval-v1 Li et al. (2023), and run it against the default reference model (text-davinci-003). Per the benchmark method, we report each model’s win rate against the reference model, across a set of 805 benign instructions. To account for model capabilities, we evaluate models against MMLU test set, which includes 14K multi-choice questions, and report the accuracy on that set. We follow the original zero-shot prompt and implementation details.666https://github.com/hendrycks/test
Appendix B Complementary Results
B.1 Characterizing GCG Suffixes – Additional Results
Fig. 9 extends GCG suffix universality analysis to Qwen2.5 and Llama3.1 models.
w/ prefilling
w/ prefilling
B.2 GCG Jailbreaks are Mechanistically Shallow – Additional Results
Fig. 10 repeats advchat’s knockout experiment while prefilling different dummy tokens, showing results consistent with the analysis in §4.1.
B.3 GCG Aggressively Hijacks the Context – Additional Results
We include additional comparisons of the dominance scores across suffix distributions, aggregated over all layers (Fig. 11), and calculated on a set of benign instructions, instead of harmful (Fig. 12).
B.4 Hijacking is Key for GCG Universality – Additional Results
Extended Hijacking Comparison, Gemma2. In Fig. 13, we extend the comparison between hijacking strength and universality, considering alternative hijacking scores (i.e., attention-based and direction-based), and making it more fine-grained.
Universality vs. Hijacking Comparison, Qwen2.5. Repeating the evaluation in §6 on Qwen2.5 (instead of Gemma2) with 100 GCG suffixes (§2), shows similar trends. In particular, inspecting the correlation between universality and hijacking per layer, Fig. 14 shows that the hijacking strength in layer 21 achieves a correlation of , -value , and confidence interval of . Focusing on layer 21 (Fig. 15), we observe that the relationship seen in Gemma2 persists, with more universal suffixes obtaining stronger hijacking.
vs. Univ Score
vs. Univ Score
vs. Univ Score
vs. Univ Score
vs. Univ Score
vs. Univ Score
B.5 Boosting GCG Universality With Hijacking Enhancement – Additional Analysis
Fig. 16 shows a fine-grained analysis of the GCG variants from §7.1, including a new variant GCG-HotInit—initializing GCG with an Adv. Injection suffix (as used for §5, and exemplified in Tab. 3), instead of the default “!!..!” initial suffix, and without modifying the objective (unlike in GCG-Hij).
Fig. 17 shows the hijacking strength throughout the GCG variants’ optimization, averaged over all the runs executed for each variant. It demonstrates that the hijacking emerges during the optimization of the suffix. Moreover, as expected, it shows GCG-Hij suffixes converge to stronger hijacking compared to GCG, and initializing the GCG optimization with strong-hijacking suffixes (GCG-HotInit) gives the optimization a head start, leading them to converge to hijacking strength between GCG and GCG-Hij. This might explain past attacks’ preference for initializing with these handcrafted jailbreak suffixes Liu et al. (2024).